OpenStudy (anonymous):

A voltage V across a resistance R generates a current I=V/R. A constant voltage of 8 volts is put across a resistance that is increasing at a rate of 0.4 ohms per second when the resistance is 5 ohms, at what rate is the current changing? ***round answer to three decimal places*** The current is changing at a rate of ________ amperes/second. please explain? Thank you!!!

OpenStudy (ash2326):

@iheartfood We have \[I=\frac{V}{R}\] Could you differentiate I with respect to T, treat V as constant and R as variable

OpenStudy (anonymous):

-V/r^2 ?

OpenStudy (ash2326):

Yes, but it's differentiated with respect to t, so you will will get \[\frac{dI}{dT}=\frac{-V}{R^2}\times \frac{dR}{dT}\]

OpenStudy (ash2326):

Do you follow?

OpenStudy (anonymous):

yes:) ....so far haha :P

OpenStudy (ash2326):

Just need to plug in the values on the left side and you get the rate of change of current

OpenStudy (ash2326):

Could you try?

OpenStudy (anonymous):

okay... :) umm so is this close (ish? haha) ? dI -8 dR --- = ----- * ------ dt 5^2 dT ? :/

OpenStudy (ash2326):

Yes, it's close "increasing at a rate of 0.4 ohms per second when the resistance is 5 ohms" When R=5 ohm \[\frac{dR}{dT}=0.4 ohm/sec\]

OpenStudy (anonymous):

ahh darn :P ohh okay so we get this? dI/dT = -8 ----- * 0.4 25

OpenStudy (ash2326):

yes Sir. It does make sense with rate of resistance increase positive, rate of current will be negative.

OpenStudy (anonymous):

ahh okay yay! :) so then we get this? -0.128 ?

OpenStudy (ash2326):

yes, correct

OpenStudy (anonymous):

ahh!! woo! thank you so much!! :D

OpenStudy (ash2326):

Welcome :)